# From Feedback to Checklists: Grounded Evaluation of AI-Generated Clinical Notes

### v1

@@ Abstract
AI-generated clinical notes are increasingly used in healthcare, but evaluating their quality remains a challenge due to high subjectivity and limited scalability of expert review. Existing automated metrics often fail to align with real-world physician preferences. To address this, we propose a pipeline that systematically distills real user feedback into structured checklists for note evaluation. These checklists are designed to be interpretable, grounded in human feedback, and enforceable by LLM-based evaluators. Using deidentified data from over 21,000 clinical encounters, prepared in accordance with the HIPAA safe harbor standard, from a deployed AI medical scribe system, we show that our feedback-derived checklist outperforms baseline approaches in our offline evaluations in coverage, diversity, and predictive power for human ratings. Extensive experiments confirm the checklist’s robustness to quality-degrading perturbations, significant alignment with clinician preferences, and practical value as an evaluation methodology. In offline research settings, the checklist can help identify notes likely to fall below our chosen quality thresholds.
@@
Karen Zhou1thanks: Work done during internship at Abridge, John Giorgi2, Pranav Mani2, Peng Xu2, Davis Liang2, Chenhao Tan1,2 1University of Chicago, 2Abridge {karenzhou, chenhao}@uchicago.edu {john, pranav, peng.xu, davis}@abridge.com
@@ 1 Introduction
To achieve these goals, we propose a method to automatically generate a checklist, given user feedback. The pipeline involves leveraging user feedback as context, prompting an LLM for candidate checklist questions, and then applying a series of refinement steps to select for the most salient questions. Finally, we conduct automatic evaluation of the final checklists, with plans for human validation and checking for generalization [^1] Our checklists are not intended for clinical deployment or for altering patient care without clinician review. An example of how a checklist would be used for scoring a note can be seen in Fig. 1.
@@
  • In our internal offline evaluation, the checklist appeared more comprehensive, diverse, and predictive of human ratings than baseline (zero-shot) checklists.
  • We demonstrate robustness of the checklist to various quality-degrading perturbations (e.g., missing salient information, poor writing flow and organization, and redundancy and hallucinations) as well as significant alignment with expert preferences.
@@ 2 Related Work
Works like Goldberg et al. (2024) use LLMs to evaluate checklist compliance, and many others have explored checklist-based evaluation with LLM-as-a-Judge for a variety of tasks (Lin et al., 2024; Que et al., 2024; Chu et al., 2025). It has been extensively demonstrated that checklists can be used to improve the reliability of LLM-evaluators and agreement with human judgments (Lee et al., 2025; Pereira et al., 2024; Cook et al., 2024; Wei et al., 2025; Li et al., 2025).
@@ 4 Pipeline & Implementation
  • We focus on the AP section of the notes, since it is a key component of clinical notes and all encounters should have this section. However, our proposed methods can be extended to other note sections.
@@
Fig. 2 shows the proposed pipeline [^2] This pipeline is illustrative, uses de-identified data, and remains subject to further validation. We will go over generation (§4.1) and refinement (§4.2) of the checklist questions below. See §5 for evaluation and results of the final checklist questions.
@@ 4.1 Checklist Generation Methods
We prompt a LLM with the instruction to generate candidate checklist questions, without any feedback or specified guidelines.
@@ 4.2 Checklist Refinement
We measure the enforceability rate for each checklist question, which is the proportion of rewritten notes that correctly fail the criteria, which is then averagied across all questions into a single enforceability score for the checklist. §D.2 further details the implementation of these unit tests. We discard questions with an enforceability score below a threshold of $0.7$.
@@
The final refinement step selects an optimal subset of checklist questions that maximizes coverage of user feedback while minimizing redundancy and checklist length. We define coverage as the proportion of feedback items addressed by the selected questions, measured using an LLM-based coverage matrix, and diversity as the average dissimilarity (1 minus Jaccard similarity) between the sets of feedback covered by each question. Then for coverage $C$, diversity $D$, weight $alpha$, and length penalty $\lambda$, we can define the following objective score:
@@
Metric Baseline ($n=10$) Feedback ($n=25$)
Coverage $\uparrow$ 0.978 0.988
Diversity $\uparrow$ 0.897 0.917
Intra-cluster Dist $\downarrow$ 0.670 0.661
Inter-cluster Dist $\uparrow$ 0.061 0.083
Table 2: Evaluation of the final checklists, with respect to feedback coverage and diversity. Diversity is additionally evaluated by intra-cluster and inter-cluster distances. The feedback checklist has better coverage and diversity than the baseline checklist.
@@ 5 Evaluation
To demonstrate our pipeline, we generate a checklist from feedback data for the AP section of the notes. Our final checklist of 25 questions is compared against a baseline checklist of 10 questions, which are written out in §E. §F.1 describes our specific generation and refinement results.
@@ 5.1 Feedback Coverage & Diversity and LLM Enforceability
Our checklist has better feedback coverage and diversity and is more LLM-enforceable than the baseline checklist, as shown in Table 2. The feedback checklist has a coverage of 0.988 and a diversity of 0.917, compared to the baseline’s coverage of 0.978 and diversity of 0.897. The feedback checklist also has a lower intra-cluster distance and a higher inter-cluster distance than the baseline checklist, indicating that it covers a wider range of feedback items while avoiding redundancy. Our feedback checklist also contains a greater quantity of LLM-enforceable questions (see §F.2).
@@ 5.2 Predictive Power
Our checklist has better predictive power than the baseline checklist as well. Our feedback checklist yields a higher accuracy of 0.70 for predicting the star ratings of the notes, compared to 0.62 for the baseline checklist. We also find that our feedback checklist has a higher macro F1 score of 0.64, compared to 0.43 for the baseline checklist. This indicates that the feedback checklist is better aligned with the star ratings of the notes.
Figure 3: The feedback checklist has a higher perturbation delta than the baseline checklist. It is more robust against perturbations, particularly for missing information, organization, and redundancy/hallucination.
@@ 5.3 Robustness against Perturbations
Our checklist is significantly more robust against perturbations than the baseline checklist. As shown in Fig. 3, the feedback checklist has an average perturbation $\Delta$ of 2.30, compared to 0.91 for the baseline checklist. The $\Delta$’s are significantly greater than 0 for all perturbations ($p<0.001$, one-sample $t$-test). In contrast, the baseline shows no significant change for the repeat_sentence and coherence_shuffle perturbations. We also confirm that the $\Delta$’s are significantly larger for our feedback checklist than the baseline, with $p<0.05$ for a two-sample $t$-test and a large effect size (Cohen’s $d=3.12$).
@@ 5.4 Correlation with Human Preference Ratings
Figure 4: Correlation with human preference ratings is significant for our checklist ($p\leq 0.05$ from a paired $t$-test), but not for the baseline checklist.
@@ 6 Conclusion
In this work, we propose a method to automatically generate a checklist for evaluating clinical notes, based on real user feedback for notes produced by a commercial Ambient AI company. We show that our generated checklists are have more desirable properties than a baseline checklist, and we define a series of evaluation metrics to assess the quality of the checklists. In particular, we show that our checklists are comprehensive, diverse, predictive of human ratings, and aligned with expert preferences. Ultimately, this work demonstrates systematic, feedback-driven evaluation for AI-generated clinical documentation.
@@
Our pipeline makes several simplifying assumptions, such as excluding "N/A" answers and cross-section comparisons, which may limit the scope and nuance of the resulting checklists. The approach is heavily reliant on LLMs for both checklist generation and evaluation, introducing potential errors due to unverified LLM outputs and a lack of ground truth labels. Additionally, checklist coverage is inherently limited by the content of user feedback, and current evaluation metrics and proxies (like internal preference ratings) have their own shortcomings. We further discuss the limitations of our approach in §G.
Future work includes scaling the pipeline to additional note sections and domains, implementing dynamic and more robust feedback filtering, and incorporating advanced evaluation methods such as feature importance analysis, human studies, and improved LLM-evaluator reasoning to further refine and validate the checklists. Extended discussion of future work can be found in §H.
@@
Our work involves the use of clinical note data and user feedback for the purpose of building automated evaluation pipelines. All proprietary data used in this work, such as user feedback, star ratings, and clinical note content, was handled in compliance with applicable data privacy policies and HIPAA regulations. All PHI was de-identified in accordance with HIPAA safe harbor standard before model training and evaluation; no direct identifiers were transmitted outside the covered environment.
We acknowledge that LLM outputs may introduce subtle errors or biases, and we flag this as an area for continued validation. Our generated checklists are designed to be interpretable and transparent, but further validation with practicing clinicians is necessary before clinical deployment.
@@ B.1 User Feedback
We retrieve ~22K encounters with user feedback of $>2$ words, non-null (1-5) star ratings. Table 1 provides some synthetic examples. The user feedback is free-text, and the star ratings are on a scale of 1-5, with 1 being the worst and 5 being the best; these are obtained from real users of an Ambient AI company that scribes clinical encounters and generates clinical notes. We specifically filter for feedback that is relevant to note quality and content, as opposed to other issues like recording or EHR integration. We use an internal tool to determine the relevant note section (e.g. HPI, A&P) for each piece of feedback and only keep the relevant feedback for each section.
@@ D.3 Optimizing for Feedback Coverage and Diversity
Since we also want to avoid overly length checklists, we also add a weight $\lambda$ for penalizing length. Finally, for a checklist of $k$ quesitons, we can define a objective score to balance coverage and diversity of a checklist, given $\alpha$ as the weight:
@@
Figure 6: Objective score (coverage and diversity) vs. length of checklist. $k=25$ is the optimal number of questions, with a good balance of coverage and diversity.
Figure 7: The enforceability score of each checklist question, indicating how LLM-enforceable each one is. Although both checklists’ average scores are the same, the feedback checklist has a greater quantity of LLM-enforceable questions.
@@ Appendix E Checklist Questions
    Are documented medical inferences or conclusions properly placed in the assessment and plan sections instead of within the HPI?
@@
    Does the assessment and plan section include a prioritized list of problems or diagnoses relevant to the visit?
@@
    Is there an accurate and specific patient management plan for each condition mentioned in the assessment and plan?
@@
    Are duplicate or redundant statements avoided in the assessment and plan?
@@
    Does the assessment and plan use third-person pronouns to ensure clarity of who will perform the actions?
@@
    Are recommendations and patient education documented clearly and specifically in the assessment and plan?
@@
    Is the assessment and plan section organized and clearly structured to facilitate future review of the patient’s management plan?
@@ F.1 Generation and Refinement
We start with the ~$7700$ feedback items that are tagged with the AP section, which are randomly split into 8 batches of approximately 900-1000 items each. A total of 97 candidate checklist questions are generated from the feedback.
We then apply the refinement steps to filter out redundant questions and tag for applicability and specificity, which results in 50 questions.
For example, the questions “Is the assessment and plan section well-organized and clearly separates different medical issues or diagnoses?” and “Is the assessment and plan organized in a way that is easy to read, with distinct separation between different patient problems?” are clustered and merged into a single question: “Is the assessment and plan section well-organized with clear separation of medical issues or diagnoses?”
These 50 questions are then evaluated for LLM enforceability, using the unit test pass rate. We find that 40 questions have a unit test threshold of at least 0.7, and we discard the remaining 10 questions. Such discarded questions include: “Is unnecessary and speculative language avoided in the assessment and plan section?” The average unit test score for the remaining questions is 0.93.
Finally, we select a subset of $k$ questions that maximize coverage and diversity. Setting $\alpha=0.5$ and $\lambda=0.0005$, we find that $k=25$ provides the optimal balance of coverage and diversity with length of the checklist (see Fig. 6).

### v2

@@ Abstract
AI-generated clinical notes are increasingly used in healthcare, but evaluating their quality remains a challenge due to high subjectivity and limited scalability of expert review. Existing automated metrics often fail to align with real-world physician preferences. To address this, we propose a pipeline that systematically distills real user feedback into structured checklists for note evaluation. These checklists are designed to be interpretable, grounded in human feedback, and enforceable by LLM-based evaluators. Using deidentified data from over 21,000 clinical encounters (prepared in accordance with the HIPAA safe harbor standard) from a deployed AI medical scribe system, we show that our feedback-derived checklist outperforms a baseline approach in our offline evaluations in coverage, diversity, and predictive power for human ratings. Extensive experiments confirm the checklist’s robustness to quality-degrading perturbations, significant alignment with clinician preferences, and practical value as an evaluation methodology. In offline research settings, our checklist offers a practical tool for flagging notes that may fall short of our defined quality standards.
@@
Karen Zhou1thanks: Work done during internship at Abridge., John Giorgi2, Pranav Mani2, Peng Xu2, Davis Liang2, Chenhao Tan1 1University of Chicago, 2Abridge {karenzhou, chenhao}@uchicago.edu {john, pranav, peng.xu, davis}@abridge.com
@@ 1 Introduction
To achieve these goals, we propose a method to automatically generate a checklist, given user feedback. The pipeline involves leveraging user feedback as context, prompting an LLM for candidate checklist questions, and then applying a series of refinement steps to select for the most salient questions. Finally, we conduct automatic evaluation of the final checklists, with plans for human validation and checking for generalization [^1] Our checklists are not intended for clinical deployment or for altering patient care without clinician review.. An example of how a checklist would be used for scoring a note can be seen in Fig. 1.
@@
  • In our internal offline evaluation, the checklist appears more comprehensive, diverse, and predictive of human ratings than a baseline (zero-shot) checklist.
  • We demonstrate the checklist’s robustness to various quality-degrading perturbations (e.g., missing information, poor writing flow and organization, and redundancy and hallucinations) as well as significant alignment with expert preferences.
@@ 2 Related Work
Works like Goldberg et al. (2024) use LLMs to evaluate compliance with a pre-existing checklist, and many others have explored checklist-based evaluation with LLM-as-a-Judge for a variety of tasks (Lin et al., 2024; Que et al., 2024; Chu et al., 2025). It has been extensively demonstrated that checklists can be used to improve the reliability of LLM-evaluators and agreement with human judgments (Lee et al., 2025; Pereira et al., 2024; Cook et al., 2024; Wei et al., 2025; Li et al., 2025).
Figure 2: Proposed end-to-end pipeline.
3 Data
@@ 4 Pipeline & Implementation
  • We focus on the Assessment and Plan (AP) section of the notes, since it is a key component of clinical notes and all encounters should have this section. However, our proposed methods can be extended to other note sections.
@@
Fig. 2 shows the proposed pipeline [^2] This pipeline is illustrative, uses de-identified data, and remains subject to further validation.. We will go over generation (§4.1) and refinement (§4.2) of the checklist questions below. See §5 for evaluation and results of the final checklist questions.
@@ 4.1 Checklist Generation Methods
We prompt a LLM with the instruction to generate candidate checklist questions, without any feedback or specified guidelines (zero-shot). This baseline is designed to isolate the contribution of clinician feedback. We chose not to adapt prior rubrics such as Stetson et al. (2012) or Burke et al. (2014) since they are Likert-style and not directly comparable to binary checklists without careful redesign, and having a clinical expert draft checklist questions from scratch has proven non-trivial and time-consuming, due to the subjectivity and exhaustiveness of the task.
@@ 4.2 Checklist Refinement
We measure the enforceability rate for each checklist question, which is the proportion of rewritten notes that correctly fail the criteria, which is then averaged across all questions into a single enforceability score for the checklist. §D.2 further details the implementation of these unit tests. We discard questions with an enforceability score below a threshold of $0.7$.
@@
The final refinement step selects an optimal subset of checklist questions that maximizes coverage of user feedback while minimizing redundancy and checklist length. We define coverage as the proportion of feedback items addressed by the selected questions, measured using an LLM-based coverage matrix, and diversity as the average dissimilarity (1 minus Jaccard similarity) between the sets of feedback covered by each question. Then for coverage $C$, diversity $D$, weight $\alpha$, and length penalty $\lambda$, we can define the following objective score:
@@
Metric Baseline ($n=10$) Feedback ($n=20$)
Coverage $\uparrow$ 0.978 0.987
Diversity $\uparrow$ 0.897 0.923
Intra-cluster Dist $\downarrow$ 0.670 0.667
Inter-cluster Dist $\uparrow$ 0.061 0.080
LLM enforceability $\uparrow$ 0.82 0.89
Table 2: Evaluation of the final checklists, with respect to feedback coverage, diversity, and LLM enforceability. Diversity is additionally evaluated by intra-cluster and inter-cluster distances. The feedback checklist beats the baseline checklist in each category.
@@ 5 Evaluation
To demonstrate our pipeline, we generate a checklist from feedback data for the AP section of the notes. From the ~$7700$ feedback items that are tagged with the AP section, 97 candidate checklist questions are generated. Following our refinement steps, we obtain a final checklist of 20 questions. Our checklist is compared against a baseline checklist of 10 questions, both of which are written out in §E. §F.1 describes our specific generation and refinement results.
@@ 5.1 Feedback Coverage & Diversity and LLM Enforceability
Our checklist has better feedback coverage and diversity and is more LLM enforceable than the baseline checklist, as shown in Table 2. The feedback checklist has a coverage of 0.987 and a diversity of 0.923, compared to the baseline’s coverage of 0.978 and diversity of 0.897. The feedback checklist also has a lower intra-cluster distance and a higher inter-cluster distance than the baseline checklist, indicating that it covers a wider range of feedback items while avoiding redundancy. Our feedback checklist also obtains a higher enforceability score, i.e., it is more answerable by LLMs (see §F.2 for the breakdown across questions).
@@ 5.2 Predictive Power
Our checklist has better predictive power than the baseline checklist as well. Our feedback checklist yields a higher accuracy of 0.70 for predicting the star ratings of the notes, compared to 0.62 for the baseline checklist. We also find that our feedback checklist has a higher macro F1 score of 0.63, compared to 0.43 for the baseline checklist. This indicates that the feedback checklist is better aligned with the star ratings of the notes.
Figure 3: The feedback checklist has a higher perturbation $\Delta$ than the baseline checklist. It is more robust against perturbations, particularly for missing information, organization, and redundancy/hallucination.
@@ 5.3 Robustness against Perturbations
Our checklist is significantly more robust against perturbations than the baseline checklist. As shown in Fig. 3, the feedback checklist has an average perturbation $\Delta$ of 2.26, compared to 0.91 for the baseline checklist. The $\Delta$’s are significantly greater than 0 for all perturbations ($p<0.001$, one-sample $t$-test). In contrast, the baseline shows no significant change for the repeat_sentence and coherence_shuffle perturbations. We also confirm that most $\Delta$’s are significantly larger for our feedback checklist than the baseline, with $p<0.05$ for a two-sample $t$-test and a large effect size (Cohen’s $d=2.94$).
@@ 5.4 Correlation with Human Preference Ratings
Figure 4: Correlation with human preference ratings is significant for our checklist ($p\leq 0.05$ from a paired $t$-test, Cohen’s $d=0.28$), but not for the baseline.
@@ 6 Conclusion
In this work, we propose a method to automatically generate a checklist for evaluating clinical notes, based on real user feedback for notes produced by a commercial Ambient AI company. We show that our generated checklists have more desirable properties than a baseline checklist, and we define a series of evaluation metrics to assess the quality of the checklists. In particular, we show that our checklists are comprehensive, diverse, predictive of human ratings, and aligned with expert preferences.
Our approach addresses a critical need in healthcare AI for scalable, clinically-grounded evaluation. By systematically distilling feedback from over 21,000 real clinical encounters, we demonstrate that user concerns can be transformed into actionable evaluation criteria. The resulting checklist serves as an interpretable, enforceable tool that can flag potentially problematic notes while maintaining transparency about evaluation criteria.
This work demonstrates the practical value of leveraging large-scale user feedback to create evaluation tools for deployed AI systems. Our pipeline provides a systematic approach to convert real user feedback into structured evaluation criteria, offering a scalable alternative to purely manual review while preserving clinical input and oversight.
Future work includes scaling the pipeline to additional note sections and domains, implementing dynamic and more robust feedback filtering, and incorporating advanced evaluation methods such as feature importance analysis, human studies, and improved LLM-evaluator reasoning to further refine and validate the checklists. Extended discussion of future work can be found in §G.
@@
Our pipeline makes several simplifying assumptions, such as excluding "N/A" answers and cross-section comparisons, which may limit the scope and nuance of the resulting checklists. Checklist coverage is inherently limited by the content of user feedback; e.g., if there are no complaints about pronoun usage, there may not be a checklist question about using correct pronouns. Checklist questions that favor comprehensive notes may penalize concise ones that are still clinically sufficient, such as in time-sensitive settings like emergency care. Our methodology could be adapted for these specialties by filtering for specialty-specific feedback, as well as for other contexts (e.g., multi-provider notes).
Our current setup treats all checklist questions with equal weight, and though this assessment is inherently subjective, we acknowledge that clinical significance varies across checklist items. Weighted scoring could be implemented automatically (e.g., Wei et al. (2025)) or added during clinical review of the rubric as a post-hoc step, without redesigning our existing methodology.
We rely heavily on LLMs in our pipeline and evaluation, which introduces its own set of challenges. When using LLMs-as-a-Judge, there is a lack of ground truth labels to determine whether clinical notes satisfy checklist questions, such as for ACI-bench notes. Exhaustive evaluation is difficult, since each checklist question can be a substantial problem on its own. As such, our current metrics are not exhaustive. Finally, we rely on proxies for positive signals, such as internal pairwise preference data, though these proxies also have their own limitations (e.g., poor agreement between raters, though we mitigate this by selecting pairs with at least two raters in agreement).
@@
Our work involves the use of clinical note data and user feedback for the purpose of building automated evaluation pipelines. All proprietary data used in this work, such as user feedback, star ratings, and clinical note content, is handled in compliance with applicable data privacy policies and HIPAA regulations. All PHI is de-identified in accordance with HIPAA safe harbor standard before model training and evaluation; no direct identifiers are transmitted outside the covered environment.
Our checklist is designed as a scalable triage tool, not a replacement for clinical judgment. Our generated checklists are designed to be interpretable and transparent, but further validation with practicing clinicians is necessary before clinical deployment. Clinicians can review the finite set of criteria, after which the checklist can be applied at scale to flag notes for expert attention. We acknowledge that LLM outputs may introduce subtle errors or biases, and we flag this as an area for continued validation. In production, the checklist will be part of a much larger evaluation suite that involves human and clinician oversight at different stages, not to be used as a single criterion for decision making.
Acknowledgments
We thank the anonymous reviewers for their comments, which helped improve this work. We also thank Zack Lipton, Elisa Ferracane-Kelley, Kaveh Moghbeli, Mike Myerburg, and the broader clinical and machine learning teams at Abridge, for their support and feedback throughout this project.
@@ B.1 User Feedback
We retrieve ~22K encounters with user feedback of >2>2 words, non-null (1-5) star ratings. Table 1 provides some synthetic examples. The user feedback is free-text, and the star ratings are on a scale of 1-5, with 1 being the worst and 5 being the best; these are obtained from real users of an Ambient AI company that scribes clinical encounters and generates clinical notes. We specifically filter for feedback that is relevant to note quality and content, as opposed to other issues like recording or EHR integration. We use an internal tool to determine the relevant note section (e.g. HPI, AP) for each piece of feedback and only keep the relevant feedback for each section.
@@ D.2 LLM Enforceability
Post-processing is done to ensure only the rewritten note is produced, i.e., no explanations. Manual verification of 50 pairs of notes spanning 27 different checklist questions has yielded a 90% validation rate (rewritten notes that correctly fail); the remaining 10% of rewritten notes are still of inferior quality (missing/inaccurate details), but do not necessarily fail the criteria of the question.
@@ D.3 Optimizing for Feedback Coverage and Diversity
Since we also want to avoid overly length checklists, we also add a weight $\lambda$ for penalizing length. Finally, for a checklist of $k$ questions, we can define a objective score to balance coverage and diversity of a checklist, given $\alpha$ as the weight:
@@
Figure 6: Objective score (coverage and diversity) vs. length of checklist. $k=20$ is the optimal number of questions.
Figure 7: The enforceability score of each checklist question, indicating how LLM enforceable each one is. The feedback checklist has both a greater quantity of perfectly enforceable questions and a higher average rate of enforceability.
@@ Appendix E Checklist Questions
    Are diagnoses and plans consistent with the patient’s reported issues and clinical findings?
@@
    Are diagnoses prioritized correctly according to the main reason for the visit?
@@ F.1 Generation and Refinement
The ~$7700$ feedback items for the AP section are randomly split into 8 batches of approximately 900-1000 items each. 97 candidate checklist questions are generated from this set of feedback.
Filtering out redundant questions and tag for applicability and specificity results in 50 questions. For example, the questions “Is the assessment and plan section well-organized and clearly separates different medical issues or diagnoses?” and “Is the assessment and plan organized in a way that is easy to read, with distinct separation between different patient problems?” are clustered and merged into a single question: “Is the assessment and plan section well-organized with clear separation of medical issues or diagnoses?”
These 50 questions are then evaluated for LLM enforceability. We find that 36 questions have a unit test threshold of at least 0.7, and we discard the remaining 14 questions. Such discarded questions include: “Is unnecessary and speculative language avoided in the assessment and plan section?” The average enforceability score for the remaining questions is 0.9.
Finally, we select a subset of $k$ questions that maximize coverage and diversity. Setting $\alpha=0.5$ and $\lambda=0.0005$, we find that $k=20$ provides the optimal balance of coverage and diversity with length of the checklist (see Fig. 6).